Dudley Bank has the following balance sheet and income statement.
Balance Sheet (in millions) Assets Liabilities and Equity Cash and due Demand from banks $ 7,500 deposits $ 15,500 Repurchase page 412
Income Statement
Interest on fees and loans $6,715
Interest on repurchase agreements 143
Interest on other investment securities 1,705
Interest on deposits in banks    60
Total interest income $8,623
Interest on deposits 3,018
Interest on debentures  1,140
Total interest expense $4,158
Net interest income $4,465
Provision for loan losses 200
Noninterest income. 950
Noninterest expenses  1,720
Income before taxes $3,495
Taxes  1,220
Net income 2275. For Dudley Bank, calculate:
Return on equity
Return on assets
Asset utilization
Equity multiplier
Profit margin
Interest expense ratio
Provision for loan loss ratio
Noninterest expense ratio
Tax ratio
Answer #1
In case of Dudley Bank
1-Return on Equity = Net Income *100 = 2275 *100 = 15.04%
Shareholder's Equity 15125
(Common Stock+Paid up Capital+Retained Earnings)
2-Return on Assets = Net Income *100 = 2275 *100 = 1.65%
Average Total Assets 137500
(Since opening total assets details are not given therefore taking closing total assets)
3-Asset Utilization = Total Sales/Interest *100 = 8623 *100 = 6.27%
(Fixed Asset Turnover Ratio) Total Assets 137500
4-Equity Multiplier = Total Assets = 137500 = 9.090909
Shareholder's Equity 15125
5-Profit Margin = Net Income *100 = 2275 *100 = 26.38%
Total Sales/Interest 8623
6-Interest Expense Ratio = Interest Expense *100 = 3018+1140 *100 = 48.22%
Total Sales/Interest 8623
7-Provison for loan loss Ratio = Provisio for loan loss *100 = 200 *100 = 2.32%
Total Sales/Interest 8623
8-Non Inerest Expense Ratio = Non Interest Expense *100 = 1720 *100 = 19.95%
Total Sales/Interest 8623
9-Tax Ratio = Tax Expense *100 = 1220 *100 = 34.91%
Taxable Income 3495